OpenStudy (ally_b):

Find the standard form of the equation of the parabola with a focus at (7, 0) and a directrix at x = -7. The options are: y = one divided by twenty eightx2 x = one divided by twenty eighty2 -28y = x2 y2 = 14x

OpenStudy (photon336):

here is my idea.  (x-h)^2 = 4p(y-k) since the vertex is in between the directrix and the focus I think that it's right at the origin. 0,0 so yeah we also know that the equation opens up sideways. so I believe that it would be in the form y^2 = x it's either between y^2 = 14x or x = 1/28 y^2